When the frequency is twice the resonant frequency, the impedance of a series RLC circuit is three times the value of the impedance at resonance. Obtain the ratios of the inductive and capacitive reactances to the resistance (XL/R and XC/R) when the frequency is twice the resonant frequency.
Answer #1

Frequency = w = 2/sqrt(LC)

Impedence = Z = 3R

XL = wL = 2*sqrt(L/C)

XC = 1/wC = 1/[2*sqrt(C/L)] = 0.5*sqrt(L/C)

Then inductive reactance to resistance = XL/R = wL/R = 2*sqrt(L/CR^2)

Then capacitive reactance to resistance = XC/R = 1/wCR = 0.5*sqrt(L/CR^2)